# Explainable Concept Generation through Vision-Language Preference Learning

### v1

@@ 3.5 The Effects of RLPO-based Concept Space Traversal
4 Experiments
Following concept-based XAI methods, we used three pre-trained classifiers, ResNet-50 [15], GoogleNet [16], and InceptionV3 [17], and two modern classifers, ViT [18] and Swin [19], as the NUT on ImageNet classes. We conducted experiments on an NVIDIA GeForce RTX 4090 GPU.
4.1 Ablation study
Table 1: Comparison between search strategies
Method Entropy ($\uparrow$) ANC ($\uparrow$) CV ($\downarrow$)
RL (Ours) 2.80 0.43 0.46
0.25 Greedy 2.40 0.21 0.96
0.5 Greedy 1.95 0.15 1.67
0.75 Greedy 1.85 0.15 1.77
An RL algorithm was used because of the large state-action space. For this study, we chose DQN as our RL algorithm because of its ability to effectively traverse through small (20 actions) and discrete action space (unique keywords). We assess the effectiveness of RL by disabling preference optimization step. As shown in Table 1, on GoogleNet, compared to $\epsilon$-greedy methods, our RL setup exhibits a higher entropy, higher average normal count (ANC) and lower coefficient of variance (CV), indicating that it is able to explore the space more broadly by picking diverse keywords.
4.2 What kind of concepts can RLPO generate?
Novel concepts. As illustrated in Fig. 2, as hypothesized, we observed that the generative model can generate new concepts that a human would not typically think of but leads activations of the DNN to trigger. To validate this hypothesis, if conducted a survey to see if humans can think of these levels of abstraction as important for the neural network to understand a certain class (Table 2). As detailed in Supplementary, we presented a random class image followed by two concepts, one generated by our method and another from a previous retrieval based method [4, 3]. We also provide options from the top priority concepts from both method and we discovered that while most participants could recognize retrieval-based concepts, only those with domain-specific knowledge could identify generated concepts. This indicates that most people can only identify concepts a small subset of what $f(.)$ learns during training.
Intuitively, when we retrieve concepts from the test class, they tend to be similar to the test images. In addition to visual observations, we verify this by computing the vector similarity of the CLIP and ResNET50 embeddings between $X_{c}$ and $X_{m}$.
Table 2: Results from the human survey. A indicates retrieval based concepts, B indicates generated concepts. Acc and Odds represent accuracy and odds, respectively.
Laymen (n=26) Expert (n=22)
Acc A 93.46% 89.55%
Acc B 8.46% 34.55%
Odds A 14.29 8.57
Odds B 0.09 0.53
Table 3: Novel concepts: $S_{c,m}$ (TCAV score), CS (Cosine similarity), ED (Euclidean distance), RCS and RED (CS and ED with ResNet50 embedding)
EAC [5] Lens [4] CRAFT [3] RLPO (Ours)
Metric C C1 C2 C3 C1 C2 C3 C1 C2 C3
$S_{c,m}(\uparrow)$ $1.0$ $1.0$ $1.0$ $1.0$ $1.0$ $1.0$ $1.0$ $1.0$ $1.0$ $1.0$
CS($\downarrow$) $0.76\pm 0.03$ $0.77\pm 0.02$ $0.72\pm 0.04$ $0.69\pm 0.05$ $0.76\pm 0.04$ $0.72\pm 0.02$ $0.73\pm 0.04$ $0.52\pm 0.04$ $0.49\pm 0.02$ $0.49\pm 0.02$
ED($\uparrow$) $7.21\pm 0.63$ $7.17\pm 0.34$ $8.02\pm 0.87$ $8.45\pm 0.96$ $7.37\pm 0.62$ $8.25\pm 0.39$ $7.98\pm 0.79$ $10.48\pm 0.5$ $10.65\pm 0.2$ $10.74\pm 0.3$
RCS($\downarrow$) $0.67\pm 0.14$ $0.5\pm 0.18$ $0.42\pm 0.1$ $0.45\pm 0.05$ $0.57\pm 0.16$ $0.5\pm 1.9$ $0.44\pm 0.07$ $0.04\pm 0.01$ $0.02\pm 0.02$ $0.03\pm 0.01$
RED($\uparrow$) $6.34\pm 2.16$ $9.7\pm 3.2$ $10.9\pm 2.8$ $11.03\pm 2.17$ $8.8\pm 3.2$ $9.9\pm 3.4$ $10.8\pm 1.9$ $16.8\pm 1.4$ $17.2\pm 0.8$ $17.6\pm 4.4$
Multiple concepts. Because the RL algorithm explores various various exlainable states, we can obtain multiples concepts with varying TCAV scores. Fig. 4 shows the top three class-level concepts identified by our method for the “zebra” class on GoogleNet. Further, the image on the left highlights where these concepts are located in the class image. This is obtained by using ClipSeg [20], a transformer-based segmentation models which can take image prompts, $X_{c}$, and highlights in a test image, $x\in X_{m}$, which part resembles the input prompt. Fig. 7 shows how our method identifies multiple concepts with importance, when compared to other methods like Anchors [21] and GradCam [22]. Table 3 highlights how concepts generated by our method are different from one extracted from the dataset.
@@ 4.2 What kind of concepts can RLPO generate?
Figure 6: Concept-shift analysis pre-post fine-tuning plot
Abstract concepts. In Fig. 2, we observe the progression of output images generated by the SD when preference optimization is applied for the initial prompt “zoo,” which was decided by the DQN, on the tiger class. These abstractions gives us hint about what the model prefers when it is looking for tiger, starting from a four-legged orange furred animal, to black and white stripes with orange furred animal, to black and white stripes with orange furred and whiskers. Though we obtain concepts with various abstractions by changing $\eta$, currently it is not possible for our method to decide $\eta$ to get a particular level of abstraction.
Complex concepts. The RL algorithm is capable of integrating multiple concepts to identify complex concept combinations. In this scenario, RL episodes conclude either when a concept, formed through sequential actions, achieves a TCAV score of 1, or when a maximum of four actions have been taken. This upper limit prevents the agent from lingering indefinitely within a single episode, ensuring there is no stagnation in the optimization of the generative model. As shown in Fig. 6, we observe that during training with multiple combination, the RL agent starts taking multiple actions initially, but as the training progresses it limits to taking one or two action combination because it starts getting good scores as over time the less number action generates better results.
4.3 Are the generated concepts reliable?
Figure 7: Comparison of concepts identified by different methods and our method.

### v2

@@ Appendix B Definition, Theorems, and Proves
Proof of Theorem 3.
At each time step $t$, two sets of samples are generated near $\mathcal{C}_{G}(t)$ using a generative function $g(.)$, denoted by $s_{1}(t)=g(\mathcal{C}_{G}(t))$ and $s_{2}(t)=g(\mathcal{C}_{G}(t))$. We define the sensitivity of these samples to the concept class using a measurable attribute, $\sigma(s)$, that quantifies the alignment or closeness of a sample $s$ to the target concept class.
The optimization step at each time step selects the sample with higher sensitivity, denoted by:
$s_{\text{opt}}(t)=\arg\max\{\sigma(s_{1}(t)),\sigma(s_{2}(t))\}$
The sample with the lower sensitivity is given by
$s_{\text{min}}(t)=\arg\min\{\sigma(s_{1}(t)),\sigma(s_{2}(t))\}$
Sample $s_{\text{opt}}(t)$ and $s_{\text{min}}(t)$ is then used to adjust $\mathcal{C}_{G}$, increasing its overall sensitivity to the concept class. Consequently, the sequence of $\mathcal{C}_{G}$ over time evolves as:
$\mathcal{C}_{G}(t+1)=(\mathcal{C}_{G}(t)\cup\ s_{\text{opt}}(t))\setminus\ s_{\text{min}}(t)$
This process incrementally increases the sensitivity of $\mathcal{C}_{G}(t+1)$ to the concept class, driven by the iterative inclusion of optimized samples.
Given that $\mathcal{C}_{N}$ is already close to the target concept class, the movement of $\mathcal{C}_{G}$ through this optimization process indirectly steers $\mathcal{C}_{G}$ towards $\mathcal{C}_{N}$. As $\mathcal{C}_{G}$ evolves in this manner, the overlap between $\mathcal{C}_{G}$ and $\mathcal{C}_{N}$ naturally increases, leading to:
$\lim_{t\to\infty}|\mathcal{C}_{G}(t)\cap\mathcal{C}_{N}|\implies\lim_{t\to\infty}\mathcal{C}_{G}(t)=\mathcal{C}_{N}.$
This results from $\mathcal{C}_{G}(t)$ containing more elements that exhibit higher sensitivity similar to those in $\mathcal{C}_{N}$, thereby increasing their intersection. ∎
B.1 Definitions
Entropy: Entropy quantifies the uncertainty or randomness inherent in a probability distribution. For a discrete random variable $X$ with possible outcomes $x_{1},x_{2},\dots,x_{n}$ and corresponding probabilities $P(X=x_{i})=p_{i}$, the entropy $H(X)$ is defined as: $H(X)=-\sum_{i=1}^{n}p_{i}\log p_{i}$, where $p_{i}$ represents the probability of outcome $x_{i}$.
Odds: Odds describe how many times an event is expected to happen compared to how many times it is not. They are often used in gambling, sports betting, and statistics. The odds of an event with probability p (where p is the probability of the event happening) are calculated as: $\frac{p}{1-p}$.
Exploration Gap (EG): quantifies the proportion of missed optimal actions, defined as $1-\text{Accuracy}$, highlighting how humans frequently miss the most optimal actions when presented with generated concepts.
Average Normalized Count (ANC): The ANC is a measure of the central tendency of the normalized action frequencies within a distribution. It provides insight into how the actions are distributed relative to the overall frequency distribution. A high ANC indicates that, on average, the action frequencies are relatively large, meaning that certain actions are more dominant. Conversely, a low ANC suggests that the actions are more evenly distributed or that lower-frequency actions have a more significant presence. Given by $\frac{1}{n}\sum_{i=1}^{n}\frac{f_{i}}{\sum_{j=1}^{n}f_{j}}$, where $f_{i}$ is the frequency of action $i$.
Inverse Coefficient of Variation (ICV): A standardized measure of concentration, calculated as the ratio of the mean to the standard deviation: $\frac{\mu}{\sigma}$. It represents how many standard deviations fit into the mean.
Feedback Cost: Feedback Cost refers to the resource(GPU) expense associated with obtaining feedback during the training of the model.
Execution Time: Execution time refers to the total time taken by a model or algorithm to complete its task from start to finish. This includes the time for data processing, model computation, and generating outputs.
Appendix C Methodology
C.1 Machine learning models we use
Neural Network Under Test ($f(.)$) : We test RLPO for all the different classification models given below.
  1. 1.
    ResNet50: We utilized a pretrained model from PyTorch torchvision pretrained models with weights initialized from ResNet50_Weights.IMAGENET1K_V2.
  2. 2.
    GoogleNet: We utilized a pretrained model from PyTorch torchvision pretrained models with weights initialized from GoogLeNet_Weights.IMAGENET1K_V1.
  3. 3.
    InceptionV3: We utilized a pretrained model from PyTorch torchvision pretrained models with weights initialized from Inception_V3_Weights.IMAGENET1K_V1.
  4. 4.
    Vision Transformer (ViT): We utilized a pretrained model from PyTorch torchvision pretrained models with weights initialized from ViT_B_16_Weights.IMAGENET1K_V1.
  5. 5.
    Swin Transformer: We utilized a pretrained model from PyTorch torchvision pretrained models with weights initialized from Swin_V2_B_Weights.IMAGENET1K_V1.
  6. 6.
    TextCNN sentiment classification model: We utilized a pretrained model from Captum library. The model was trained on IMBD sentiment dataset.
TCAV logistic model : We utilized a logistic regression model to address classification tasks in TCAV instead of the default SGD (Stochastic Gradient Descent) classifier. This decision was based on our observation that the SGD classifier produced high variance TCAV (Testing with Concept Activation Vectors) scores, which indicated inconsistent model behavior across different runs. We configured the model to perform a maximum of 1000 iterations (max_iter=1000).
Stable Diffusion v1-5 with LoRA : We used our base generation model as SD v1-5 and updated its weights using LoRA during preference optimization step. This version of SD was finetuned from SD v1-2 on "laion-aesthetics v2 5+" dataset with 10% drop in text-conditioning for better CFG sampling. In our experiments, we kept LoRA rank to 8 with a scaling factor of 8 and initial weights were defined from a gaussian distribution. We only targeted the transformer modules of U-Net in the SD architecture.
DQN : We use a DQN with specific parameters tailored to effectively navigate a vast search space. We utilized a small buffer size of 100, which limits the number of past experiences the model can learn from, encouraging more frequent updates. The exploration rate was set at 0.95, prioritizing exploration significantly to ensure thorough coverage of the search space. The batch size was configured to 32. We set the discount factor to 0.99 and the update frequency was set at every four steps. The model updates its parameters with a the soft update coefficient of 1.0. Gradient steps was set to 1 indicating a single learning update from each batch, and gradient clipping was capped at 10 to prevent overly large updates.
BLIP: We utilize the Bootstrapped Language Image Pretraining (BLIP) model for the task of Visual Question Answering (VQA). This model, sourced from the pre-trained version available at ’Salesforce/blip-vqa-capfilt-large’, is designed to generate context-aware responses to visual input by leveraging both image and language understanding. The large variant of the BLIP model is fine-tuned for VQA, allowing it to effectively interpret and answer questions based on the visual content provided.
C.2 RLPO algorithm
Algorithm 2 shows the complete algorithm of the algorithm shown in Section 3 algorithm 1.
Algorithm 2 DQN Algorithm with DPO and Adaptive Reward
1:Input: Set of test images, $f(.)$
2:Initialize Q-network $Q_{\theta}(s,a)$ with random weights $\theta$
3:Initialize replay buffer $\mathcal{D}$ and adaptive parameter $\xi\leftarrow 0.1$
4:for each episode do
5:     for each time step $t$ do
6:         Observe state $s_{t}$ and select action $a_{t}$ based on $Q$ ($\epsilon$-greedy)
7:         Execute $a_{t}$ and generate 10 images, divided into two groups $G_{1}$ and $G_{2}$
8:         Evaluate TCAV scores $\overline{TCAV}_{1}$ and $\overline{TCAV}_{2}$
9:         if  $\max(\overline{TCAV}_{1},\overline{TCAV}_{2})\leq 0.7$  then
10:              Update policy to favor higher $\overline{TCAV}$ group and perform DPO
11:              Update $\xi\leftarrow\min(1,\xi+\text{increment})$
12:         else
13:              Set $\xi\leftarrow 1$
14:         end if
15:         Compute reward $r_{t}=\xi\cdot\max(\overline{TCAV}_{1},\overline{TCAV}_{2})$
16:         Store transition $(s_{t},a_{t},r_{t},s_{t+1})$ in $\mathcal{D}$
17:         Sample a mini-batch from $\mathcal{D}$
18:         for each sampled transition $(s_{i},a_{i},r_{i},s_{i+1})$ do
19:              Compute target $y_{i}=r_{i}+\gamma\max_{a^{\prime}}Q_{\theta^{\prime}}(s_{i+1},a^{\prime})$
20:         end for
21:         Compute loss $L(\theta)=\frac{1}{N}\sum_{i=1}^{N}(y_{i}-Q_{\theta}(s_{i},a_{i}))^{2}$
22:         Perform a gradient descent step to update $\theta$
23:         Periodically update target network: $\theta^{\prime}\leftarrow\tau\theta+(1-\tau)\theta^{\prime}$
24:     end for
25:end for
26:Output: Set of concept images
C.3 Preprocessing for generating the action space
Steps not discussed in Section 3.2.
Figure 11: Seed prompt pipeline
Each patch from the test images is passed to the VQA model to extract relevant and useful information about the corresponding class. In this study, we choose BLIP Li et al. (2022) as our VAQ model. We posed a set of targeted questions to the VQA model, aiming to gain insights into the class-specific features represented in the patches. The questions are designed to probe various aspects of the image patches, helping the model focus on class-defining attributes.
  1. 1.
    “What is the pattern in the image?”
  2. 2.
    “What are the colors in the image?”
  3. 3.
    “What is the background color of the image?”
  4. 4.
    “What is in the background of the image?”
  5. 5.
    “What is the primary texture in the image?”
  6. 6.
    “What is the secondary texture in the image?”
  7. 7.
    “What is the shape of the image?”
We then remove stop words and duplicates from the generated responses using lemmantizing and perform a cross-similarity check using CLIP between all the unique words and further filtered words which are more than 95% similar. To further select most relevant keywords to the class images, we perform a VLM check using class images and the extracted keyword to get the softmax score of how much the keyword and image are related. This score is then averaged over all the class images and this average is use to sort the keywords. Now, from the sorted keywords, we select top 20 keywords as our RL action space. The cross-similarity and VLM check are inspired from Zang et al. (2024) where they used a similar filtering setup to remove potentially useless concepts.
C.4 Exploration of Random Gibberish Prompts as Seed Prompts
In this experiment we don’t use a VQA to get seed prompts. We choose a random list of incoherent prompts, example shown in Fig. 12. We found that for these prompts it takes a really long time to get some meaningful explanation and in most cases lead to random generation, thus showing importance of starting from proxy concepts.
Gibberish Seed Prompts 1. dKgN MTW8bvbxB6aW1L2TfTuTYZK3He0urbEEmclEpY 2. se-L8fPe19ZzUmuM uDYVYusFnYtNZeFM1YqXdE57Y7OMD3Z80cKwLo5 3. CzKLTlZZnWHjtBn80wIfC z8O 4. mhtxqyH2FBEC 5. SWEC6Wlqfpqaz PQjoGrxIuzm m2ua8oGJySIeG2NqCG9BBvU9Eerj7wheWk7j-t
Figure 12: Sample gibberish seed prompt used with RLPO on GoogleNet classifier to generate concepts.
C.5 Preference optimization update for state space
Steps not discussed in Section 3.4.
The candidate concepts serve as the initial states for the RL agent. From these initial states, the agent takes actions $a\in\text{Keywords}$ that leads to multiple subsequent possible states using $g(.)$. These states are then grouped, and the group’s sensitivity is compared against INUT using TCAV scores. A higher TCAV score suggests higher sensitivity, indicating that the group is more aligned with INUT.
We employ preference optimization over the grouped states to guide states towards INUT. To prevent the model from skipping over explainable states and directly reaching INUT, we introduce a threshold that limits the application of preference optimization at each step as shown in equation 17.
Given two groups of samples $\text{Given two groups of samples }G_{1}\text{ and }G_{2}\text{ with their average TCAV scores }\overline{TCAV}_{1}\text{ and }\overline{TCAV}_{2}:$
if $\text{if }\max(\overline{TCAV}_{1},\overline{TCAV}_{2})\leq 0.7,\text{ update }\pi\text{ to favor the group with higher }\overline{TCAV}.$ (17)
To optimize $g(.)$ to find better proxies, for each step in the environment we utilized average TCAV scores $\overline{TCAV}_{1}$ and $\overline{TCAV}_{2}$ from $G_{1}$ and $G_{2}$ to decide between preferred and unpreferred concepts. Lets say $\overline{TCAV}_{1}\succ\overline{TCAV}_{2}$, than we optimize $g(.)$ over the sample S defined as $S=\{(a,x^{g1}_{0},x^{g2}_{0})\}$, where $x^{g1}_{0}$ and $x^{g2}_{0}$ are the sample points from the groups on action $a$. We optimize $g(.)$ using objective 18 to get a new optimzed $g^{\prime}(.)$  Wallace et al. (2023).
$L(\theta)=$ $-\mathbb{E}_{(x^{g1}_{0},x^{g2}_{0})\sim S,t\sim U(0,T),x^{g1}_{t}\sim q(x^{g1}_{t}|x^{g1}_{0}),x^{g2}_{t}\sim q(x^{g2}_{t}|x^{g2}_{0})}\log\sigma\left(-\beta T\omega(\lambda_{t})\right.$
$\left.\left(\|\epsilon^{G1}-\epsilon_{g^{\prime}(.)}(x^{G1}_{t},t)\|_{2}^{2}-\|\epsilon^{G1}-\epsilon_{g(.)}(x^{G1}_{t},t)\|_{2}^{2}\right.\right.$
$\left.\left.-\left(\|\epsilon^{G2}-\epsilon_{g^{\prime}(.)}(x^{G2}_{t},t)\|_{2}^{2}-\|\epsilon^{G2}-\epsilon_{g(.)}(x^{G2}_{t},t)\|_{2}^{2}\right)\right)\right)$ (18)
where $x^{*}_{t}=\alpha_{t}x^{*}_{0}+\sigma_{t}\epsilon^{*}$, $\epsilon^{*}\sim\mathcal{N}(0,I)$ is drawn from $q(x^{*}_{t}|x^{*}_{0})$. $\lambda_{t}={\alpha^{2}_{t}}/{\sigma^{2}_{t}}$ is the signal-to-noise ratio, and $\omega(\lambda_{t})$ is weighting function (constant in practice).
C.6 TCAV setting for different NUT models
We tested NUT models on different layers and classes and the summary of our setting across different NUT models is described in table 5.
Table 5: TCAV setting across different NUT models
Models Layers ImageNet Classes
ResNet50 penultimate layer Jay, Tiger, Rabbit & Zebra
GoogleNet inception4e layer Goldfish, Tiger, Zebra & Police Van
InceptionV3 Mixed_7c layer Goldfish, Tiger, Lionfish & Basketball
Vision Transformer (ViT) heads layer Goldfish, Golden Retriever, Tiger & Cab
Swin Transformer head layer Goldfish, Jay, Siberian husky & Tiger
Appendix D Experiments
D.1 Computing resources
The experiments were conducted on a system equipped with an NVIDIA GeForce RTX 4090 GPU, 24.56 GB of memory, and running CUDA 12.2. The system also featured a 13th Gen Intel Core i9-13900KF CPU with 32 logical CPUs and 24 cores, supported by 64 GB of RAM. This setup is optimized for high-throughput computational tasks but the experiments are compatible with lower-specification systems.
D.2 Human and LLM-based AI Feedback Mechanisms
We test other feedback mechanism in RL by replacing the XAI-TCAV feedback with AI and Human feedback’s. Herer we discuss the experiental setup and configuration for both experiements.
AI feedback: GPT-4 is leveraged to evaluate the explanatory power of image sets by focusing on concepts related to a target class, a method that aligns with the growing trend of incorporating AI-driven feedback Bai et al. (2022). This approach involves sending a structured prompt to an LLM, asking it to score how well two sets of images explain a target class using a specified concept. The process involves the following.
  1. 1.
    Image Encoding: The images from two sets (concept1 and concept2) are first converted into a base64 format to ensure they can be transmitted via the request as encoded strings.
  2. 2.
    Structured Prompt: A detailed and specific prompt is crafted for the LLM. It asks the model to assess the quality of explanation each image set provides for a particular class through the lens of a specific concept. The prompt used is “Please evaluate each of the following sets of images for how well they explain the class {class_name} via the concept {concept_name}. For each set, provide a numerical score between 0 and 1 (to two decimal places)” The prompt clearly defines how the model should respond, asking for a numerical score between 0 and 1, where:
    1. (a)
      0 indicates that the image set does not explain the class at all via the concept.
    2. (b)
      1 indicates that the image set perfectly explains the class via the concept.
  3. 3.
    LLM-Based Scoring: Once the prompt is sent to the LLM, it evaluates the image sets and provides scores based on its learned knowledge and understanding. The response is parsed to extract the scores for each set of images.
Human feedback: In this experiment, 7 computer science majors provided live feedback after each step of a reinforcement learning process, leveraging their prior knowledge of reinforcement learning with human feedback (RLHF) mechanisms Christiano et al. (2017). The feedback from all participants was averaged to serve as the reward for each step in the RL process. Given the abstract nature of the initial concepts, participants needed to take time to thoughtfully assess each step, which contributed to a lengthier feedback cycle.
D.3 Additional results and analysis
To validate our method for its ability to generate concepts, we tested it with different models and classes. We started it on traditional models, GoogleNet and InceptionV3, and then extended it to transformer-based models, Vision Transformer (ViT) and Swin Transformer, pre-trained on ILSRVC2012 data set (ImageNet) Krizhevsky et al. (2017). We show additional plot in various classes shown in Fig 13,14,15,16,17.
Figure 13: Explanation plot of Cab classification by ViT from RLPO.
Figure 14: Explanation plot of Basketball classification by InceptionV3 from RLPO.
Figure 15: Explanation plot of Lionfish classification by InceptionV3 from RLPO.
Figure 16: Explanation plot of Golden Retriever classification by ViT from RLPO.
Figure 17: Explanation plot of Tiger classification by GoogleNet from RLPO.
D.3.1 Cumulative rewards
The cumulative rewards during training for GoogleNet and InceptionV3 is shown in Fig. 18. For ViT and Swin Transformer it is shown in Fig. 19. This figure illustrates the steady accumulation of rewards over time as they interact with the reinforcement learning environment. All models demonstrate a steady increase in cumulative rewards, the classes with higher reward peak reaches its explinable state faster.
Figure 18: Cumulative rewards on traditional models.
Figure 19: Cumulative rewards on transformer models.
D.3.2 Action Selection Optimization During RLPO Training
@@
As shown in Fig. 20, during training with multiple combinations of seed prompts, we observe that the RL agent initially explores various action combinations. However, as training progresses, individual actions become more optimized due to preference optimization (PO). This leads the agent to prefer fewer action combinations, since just choosing one or two actions makes the agent reach an explainable state.
D.3.3 Concept heatmap
To determine the relationship between generated concepts and test images, we made use of CLIPSeg transformer model (Lüddecke & Ecker, 2022). We passed generated concepts as visual prompts and test images as query images into the model and it returns a pixel-level heatmap of the probability of visual prompt in the query image. Fig. 2122 showcases some examples on concept heatmap indicating the presence of the concept in the image.
Figure 21: Van class with “white blue and yellow”, Lion fish class with “zebra” seed prompt.
Figure 22: Basketball class with “basket” seed prompt, Tiger class with “orange black and white” seed prompt.
D.3.4 C-deletion
The central idea behind c-deletion in explainability is to identify and remove parts of the input context that are not crucial for the decision-making process, allowing for clearer insights into how the model arrives at its predictions or actions.
Figure 23: The figure shows c-deletion taking place for different images from “tiger” class over time for “orange black and white” seed concept.
Figure 24: The figure shows c-deletion taking place for different images from “cab” class over time for “yellow and white” seed concept.